OpenStudy (anonymous):

What is the perimeter of a triangle with the given side lengths? s1 = 2x + 1 cm s2 = 3x + 2 cm s3 = x + 3 cm

OpenStudy (anonymous):

P =(box) x + (box) cm

OpenStudy (ahsome):

TO get the perimeter, @thatoneman, simply add all the sides together

OpenStudy (jmark):

Perimeter P=s1+s2+s3 s1 = 2x + 1 cm s2 = 3x + 2 cm s3 = x + 3 cm P=2x+1+3x+2+x+3 P=6x+6 P=6(x+1) Follow us at : https://plus.google.com/+EduriteScoreMore/posts